# Diverging Towards Hallucination: Detection of Failures in Vision-Language Models via Multi-token Aggregation

### v1

@@ 5.1 Probabilistic Model
be the sequence of decoder-side embeddings (i.e for this setting, logits corresponding to each output token), and let $Y\!\in\!\{0,1\}$ denote the (unknown) ground-truth reliability label ($Y=1$: truthful, $Y=0$: hallucinated). We assume a non-informative prior $\Pr(Y\!=1)=\Pr(Y\!=0)=\tfrac{1}{2}$. A pretrained reliability head $f_{\theta}:\mathbb{R}^{d}\!\to(0,1)$ maps each token embedding to a Bernoulli parameter
$p_{\ell}\;=\;f_{\theta}(x_{\ell})\;=\;\Pr(Y\!=1\mid x_{\ell}),\qquad 0\leq\ell<T.$
Conditional on $Y$, the evidence variables $\{x_{\ell}\}$ are taken to be independent—the classical assumption behind neural likelihood-ratio tests, which we adopt primarily for improvement of tractability Cranmer et al. (2016). The training is handled through the common objective of minimization of cross entropy loss, as illustrated in Appendix B.1.
5.2 Multi-Token Log-Likelihood Ratio
For every prefix length $k\!\in\!\{1,\dots,K\}$ (with $K\!\leq\!10$ in experiments) we compute the masked log likelihood under each hypothesis:
$\ell^{(k)}_{1}$ $=\sum_{\ell=1}^{k}\,\log p_{\ell},$ $\ell^{(k)}_{0}$ $=\sum_{\ell=1}^{k}\,\log(1-p_{\ell}).$ (5)
Their difference is the cumulative log-likelihood ratio (LLR):
$\boxed{\,\Lambda^{(k)}=\ell^{(k)}_{1}-\ell^{(k)}_{0}=\sum_{\ell=1}^{k}\log\!\frac{p_{\ell}}{1-p_{\ell}}\,}.$
Equation (1) is computed in the validation loop by summing $\log p_{\ell}$ and $\log(1-p_{\ell})$ across locations and subtracting the two running totals. In practice, we discuss handling ragged batches of uneven sentence length in Appendix B.3. With equal priors, the maximum-a-posteriori (MAP) decision reduces to a sign test on the LLR:
$\hat{Y}^{(k)}\;=\;\begin{cases}1&\text{if }\Lambda^{(k)}\geq 0,\\[3.0pt] 0&\text{otherwise}.\end{cases}$
We note that utilization of a sign test can be interpreted as a threshold of $\tau\!=\!0$. We note that we leave exploration of other values of $\tau$ to improve performance for future work.
We empirically validate MTRE through extensive experiments (see Section 6), demonstrating substantial performance improvements over existing single-token baselines across multiple challenging tasks.
6 Experiments and Results
To empirically validate our method experiments are conducted with the MathVista Lu et al. (2023), MM-Safety-Bench Liu et al. (2023a), MAD-Bench Li et al. (2023), and four separate arithmetic and counting tasks from Rahmanzadehgervi et al. (2024) (see Appendix A for more details on all datasets). We test on outputs produced by open-source models including LLaVA-v1.5 (7B) Liu et al. (2023b), mPLUG-Owl Ye et al. (2023), LLaMA-Adapter (V2) Gao et al. (2023), and MiniGPT-4 Zhu et al. (2023), and unless explicitly specified, we use the 7B version of the models. We denote all prompts in Appendix D. All tested VLM’s have been reported to readily generate undesirable content under certain instructions.
6.1 Baselines
First Token Linear Probing Zhao et al. (2025)
Proposed by Zhao et al. (2025), the first-token linear probing technique serves as a baseline for comparison with our method. Linear probing evaluates whether specific information can be linearly extracted from representations learned by a model. Given a representation vector $\mathbf{h}\in\mathbb{R}^{d}$ (e.g., the logits corresponding to an output token), linear probing involves training a simple linear classifier, typically logistic regression for binary tasks, to predict a label $y\in\{0,1\}$.
$z=\mathbf{w}^{\top}\mathbf{h}+b$
$\hat{y}=\sigma(z)=\frac{1}{1+e^{-z}}$
P(True)Kadavath et al. (2022a); Steyvers et al. (2025)
Initially developed for uni-modal large language models Kadavath et al. (2022a), P(True) is a self evaluation technique to determine if an answer is: A) True or B) False, we extend this approach by applying it to open source vision-language models. For the LLM setting, the authors utilize the raw probability that a model assigns to the proposition that a given sample is the correct answer to a question. To achieve this, the authors first design a prompt, for example:
Question: Who was the first president of the United States?
Proposed Answer: George Washington
Is the proposed answer:
(A) True
(B) False
The proposed answer is:
Image:<Image Here>
Question: Who was the first president of the United States?
For comparison, we utilize the Youden index optimal cutoff point Fluss et al. (2005) obtained from the training set scores to obtain accuracy and F1 score on the validation set.
6.2 Metrics
We evaluate our methods primarily on VLM responses generated from Type 1 and Type 2 questions. Type 1 questions are questions that simply ask the VLM to predict the correct mapping of the original question in the benchmark. For Type 2 questions, the VLM is asked to analyze its Type 1 outputs and determine if its solution was correct. Additionally, we take note of some of the practical desiderata discussed in Zhao et al. (2025) and utilize different prompts: (1) OE, asking the original open-ended questions in the benchmark; (2) OEH, asking each question with a hint indicating the question may be unanswerable / may generate harmful content / may be deceptive; (3) MQ, asking a meta-question like “Is this question answerable or not?”. We note that asking the VLM different types of questions is primarily to extract different sets of outputs to evaluate the performance of each method. Each method will be using the outputs generated by either Type 1 or Type 2 as input for all tasks.
We assess the performance of MTRE using accuracy, F1 score, and the area under the receiver operating characteristic curve (AUROC) for all tasks utilizing Type 1 and Type 2 questions. We demonstrate the advantages of our method by comparing it to linear probing and P(True), with all metrics computed relative to the ground-truth labels for each evaluation type.
6.3 Discussion
We first empirically verify that our method can succeed on tasks where the first logit alone is sufficient to achieve strong performance. We begin by showcasing the final classification performance on Li et al. (2023), and Liu et al. (2023a) using Type 1 responses. Next, we discuss a potentially difficult problem of detecting correct self-evaluation. We then proceed to investigate the discriminative potential of outputs generated by arithmetic-centric tasks Lu et al. (2023); Rahmanzadehgervi et al. (2024) which are known to be particularly challenging for VLMs to solve and often lead to an increased rate of hallucinated outputs.
Final Classification Task Performance
We begin with reproducing the case proposed by Zhao et al. (2025), and use the VLM output logits from the Type 1 settings across 24 different configurations (3 Prompts x 4 VLMs x 2 Datasets) to evaluate Linear Probing, P(True), and our proposed MTRE method—on their ability to improve the final classification performance of the open-source models on the original benchmark tasks.
Table 1: Averaged performance on final classification task for each method across all models utilizing Type 1 responses as input (for exact measurements on all 24 configurations, see Appendix C). We show that our method is still able to perform similarly to a model optimized to evaluate on the first token. We utilize the VLM’s base performance as the baseline.
Section Method MM-Safety-Bench MAD-Bench
Acc Auc F1 Acc Auc F1
OE Linear Probing 89.94 96.44 90.31 90.57 96.08 90.46
P(True) 63.18 65.21 70.21 63.50 68.25 65.39
MTRE 89.67 96.16 89.26 89.46 95.17 89.01
Baseline 64.14 74.50 62.31 55.46 66.73 55.46
OEH Linear Probing 87.45 94.89 87.59 89.54 95.63 89.31
P(True) 62.31 64.86 63.86 65.88 70.43 62.57
MTRE 88.32 95.18 88.78 86.61 93.74 86.53
Baseline 66.22 70.98 65.55 59.24 70.60 59.24
MQ Linear Probing 88.43 95.04 88.68 88.73 95.28 88.50
P(True) 55.65 59.10 57.54 62.44 65.91 63.86
MTRE 88.91 95.69 89.23 88.71 94.81 88.85
Baseline 49.29 28.98 53.44 62.67 72.09 81.11
Table 1 shows that Linear Probing and our method are capable of using the logits as embeddings to correctly identify the correct classification to the benchmark task. We find that this aligns with our findings 2, as we see that Linear Probing can leverage the early divergence between samples.
Self Evaluation Task Performance
We construct a more potentially difficult task of determining if the VLM’s Type 2 assessment is correct. For a VLM response to be correct entails either confidently reaffirming a valid original answer or acknowledging its inaccuracy.
Table 2: Averaged self-evaluation performance for each method across all models utilizing Type 2 responses as input (for exact measurements on all 24 configurations, see Table 6) in the Appendix.
Section Method MM-Safety-Bench MAD-Bench
Acc Auc F1 Acc Auc F1
OE Linear Probing 54.85 52.53 64.78 77.23 78.57 80.41
P(True) 48.09 49.41 44.09 60.07 59.05 54.45
MTRE 69.05 59.30 79.49 81.31 82.40 84.37
OEH Linear Probing 48.04 53.12 39.10 74.86 83.02 79.16
P(True) 58.24 57.31 51.99 55.80 57.78 56.75
MTRE 64.19 65.65 61.99 79.18 86.96 81.09
MQ Linear Probing 59.07 61.16 53.94 80.08 86.41 81.42
P(True) 51.14 52.62 48.51 58.72 59.28 63.63
MTRE 68.34 70.53 63.00 82.41 87.94 82.89
We again test with the same 24 configurations used previously in the Type 1 final classification task setting. We begin to observe the first indications that additional logits may be beneficial. Specifically, on the MM-Safety-Bench (results shown in Table 2), and MathVista dataset (results shown in Table 3), relying solely on the first logit of Type 2 responses results in diminished performance compared to our method, which incorporates multiple logits in its prediction. Table 3 suggests that our results on MathVista improve the performance of the original work done by roughly $5.8\%$, suggesting that there indeed is more discriminative power in the logits generated from MathVista.
Table 3: Performance of various models on MathVista Self Evaluation tasks using Type 2 responses.
Model Method Accuracy AUC F1 Score
LLAVA-7B Linear Probing 67.5 74.31 58.51
P(True) 56.6 62.22 48.79
MTRE 74.1 80.80 64.17
LLAMA-Adapter Linear Probing 70.7 73.50 57.93
P(True) 69.0 68.52 43.77
MTRE 76.6 79.25 63.28
MPLUG-Owl Linear Probing 67.7 75.11 66.50
P(True) 50.9 25.25 0.00
MTRE 73.2 79.70 70.36
MiniGPT-4 Linear Probing 67.3 71.52 54.37
P(True) 36.4 35.52 53.16
MTRE 72.5 76.15 59.20
Table 4: Performance by method on each arithmetic dataset in the Hallucination Detection Task using Type 1 responses. Full experiment of all arithmetic hallucination tasks can be found in Table 7.
(a) Overlapping Circles
Method Acc Auc F1
Lin. Prb. 81.71 87.06 77.88
P(True) 61.77 68.88 58.69
MTRE 88.53 92.44 85.80
(b) Overlapping Triangles
Method Acc Auc F1
Lin. Prb. 85.18 84.53 82.27
P(True) 76.48 70.04 75.51
MTRE 90.86 91.33 91.05
(c) Intersecting Lines
Method Acc Auc F1
Lin. Prb. 84.07 88.80 70.02
P(True) 56.67 57.70 41.65
MTRE 86.58 89.55 72.41
(d) MathVista
Method Acc Auc F1
Linear Probe 73.50 71.43 40.99
P(True) 69.75 55.72 23.73
MTRE 75.21 71.31 36.72
(e) Nested Squares
Method Acc Auc F1
Linear Probe 64.34 52.67 42.90
P(True) 54.79 56.87 41.10
MTRE 88.83 91.59 82.78
All metrics are averaged over four VLMs (LLAVA-7B, LLAMA-Adapter, mPLUG-Owl, MiniGPT-4).
Determining Model Hallucination on Arithmetic Exercises
Next, we formulate a task aimed at determining whether the VLM’s initial answer of a Type 1 question was likely correct. To this end, we employ logical datasets with concrete, verifiable answers—datasets that have proven particularly challenging for VLMs to handle accurately. We show that indeed for counting tasks (One, Two, Three) the initial logits for hallucinated samples rarely diverge from the non-hallucinated samples for these datasets, and are difficult for any of the first logit methods to distinguish if the initial answer was correct. We display our findings in Table 4. We find that the first logit continues to lose the performance in determining if a VLM has hallucinated.

### v2

@@ 4.1 Token Level Training for Reliability Classifier
be the sequence of $T_{i}$ decoder-side embeddings (i.e., for this setting, the logits $\ell$ corresponding to each output token), and let $Y_{s_{i}}\in\{0,1\}$ denote the binary ground-truth reliability label ($Y_{s_{i}}=1$: truthful, $Y_{s_{i}}=0$: hallucinated) of sentence $s_{i}$. We then construct a token-level dataset by assigning each decoder-side embedding $x^{s_{i}}_{t}$ the corresponding label $Y_{s_{i}}$ inherited from its origin sentence $s_{i}$
$\mathcal{D}=\{(x^{s_{i}}_{t},Y_{s_{i}})\;\mid\;i=1,\dots,|\mathbb{S}|,\;t=1,\dots,T_{i}\}.$
Once the dataset has been shuffled with respect to its origin test or training set, we train a reliability classifier $f_{\theta}$ to predict $p_{j}$ on dataset $\mathcal{D}$ with a regularized binary cross-entropy objective:
$\mathcal{L}(\theta)=-\frac{1}{|\mathcal{D}|}\sum_{(x,Y)\in D}\Big[Y\log f_{\theta}(x)+(1-Y)\log(1-f_{\theta}(x))\Big]\;+\;\lambda\lVert\theta\rVert_{2}^{\,2}.$
In our implementation, the reliability classifier $f_{\theta}$ is chosen to be an attention-based neural network that projects the input features into a shared embedding space, applies multiple stacked multi-head self-attention layers to capture feature dependencies, aggregates the contextualized representations through adaptive average pooling, and finally passes them through fully connected layers with nonlinearities and dropout before producing a scalar reliability score via a sigmoid activation.
4.2 Token Level Aggregation For Sentence Classification
Given the reliability classifier $f_{\theta}$, we can compute the token-level reliability of each token on a generated response $s_{i}$:
$p^{s_{i}}_{t}\;=\;f_{\theta}(x^{s_{i}}_{t})\in[0,1],\ t=1,\dots,T_{i},$ (3)
where $p^{s_{i}}_{t}$ is a per-token proxy approximating the reliability $\Pr(Y_{s_{i}}{=}1\mid x^{s_{i}}_{t})$ of the logit $x^{s_{i}}_{t}$.
Then, MTRE computes the per-token Log Likelihood Ratio (LLR) $z_{s_{i},t}$ and aggregates it into a statistic (Algorithm 1), which we refer to as the evidence of the generated sentence:
$L_{s_{i},\tau_{s_{i}}}=\sum_{t=0}^{\tau_{s_{i}}}z_{s_{i},t}=\sum_{t=0}^{\tau_{s_{i}}}\log\!\frac{p^{s_{i}}_{t}}{1-p^{s_{i}}_{t}}.$ (4)
where the evidence length $\tau_{s_{i}}\leq T_{i}$ is a tunable parameter controlling how long evidence is accumulated. Intuitively, the evidence $L_{s_{i},\tau_{s_{i}}}$ quantifies the cumulative support for $s_{i}$ being reliable versus hallucinated, in the spirit of sequential probability ratio tests (Wald, 1992).
Consequently, the maximum-a-posteriori (MAP) decision rule for MTRE reduces to:
$\hat{Y}_{s_{i}}\;=\;\begin{cases}1&\text{if }L_{s_{i},\tau_{i}}\geq\delta,\\[3.0pt] 0&\text{otherwise},\end{cases}$
where $\delta=0$ corresponds to the equal-prior assumption. In the next subsection 4.3, we describe how the training dataset is used to calibrate the evidence length $\tau_{s_{i}}$—resulting in a variant of MTRE called MTRE-$\tau$—and to account for unequal class priors via out-of-fold calibration.
4.3 Parameter Calibration via Cross-Fitting
While setting the evidence length $\tau_{s_{i}}$ can be done via domain knowledge, in this subsection we present a variant of MTRE: Multi-Token Reliability Estimation $\bm{\tau}$ (MTRE-$\bm{\tau}$), a procedure which uses the training dataset to estimate $\tau_{s_{i}}$ and account for a non-uniform prior via out-of-fold (OOF) training calibration.
We describe the procedure of MTRE-$\tau$ through four distinctive steps:
  1. 1.
    Cross-fit OOF score collection. MTRE-$\tau$ begins with partitioning the training subset $\mathbb{S}_{\text{train}}$ into $K_{\text{cv}}$ stratified folds, primarily used to collect empirical estimates on how a trained reliability model may behave when given unseen data with respect to the training data. Explicitly, for each fold $j$, the MTRE-$\tau$ procedure:
    • Trains the reliability head $f_{\theta}$ on $\mathbb{S}_{\text{train}}\setminus\text{fold}_{j}$ using the Token-Level Training Algorithm 3.
    • Collects LLR $z_{s_{i},t}$ and corresponding ground truth labels $Y_{i}$ pairs ($z_{s_{i},t}$, $Y_{i}$) from $\text{fold}_{j}$ using Collection Algorithm 2 with $\tau_{s_{i}}=T_{i}$ (where $T_{i}$ is the length of $s_{i}$ or a user defined max).
  2. 2.
    Prior estimation. After OOF LLR pairs ($z_{s_{i},t}$, $Y_{i}$) have been collected from $K_{cv}$ folds, to handle non-equal-prior, MTRE-$\tau$ trains a learnable scalar $C>0$ that shifts the decision boundary, replacing the fixed threshold with a data-driven dynamic threshold. We estimate $C$ from OOF log-likelihood ratio statistics by minimizing token-broadcasted binary cross-entropy:
    $C^{\star}\in\arg\min_{C>0}\;\frac{1}{\sum_{i=1}^{N}T_{i}}\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}\mathrm{BCE}\!\left(\sigma\!\Big(\tfrac{z_{s_{i},t}}{C}\Big),\,Y_{i}\right).$ (4)
    The calibrated scores $z^{c}_{i,t}=z_{i,t}/C^{\star}$ thus correspond to a MAP test with a learned prior, allowing MTRE to adapt its threshold dynamically across datasets.
  3. 3.
    Evidence Length $\tau$ estimation. Once token LLR has been trained, the goal of this step is to estimate the termination token for each sentence $\tau_{s_{i}}$ such that the evidence aggregation halts when $L_{s_{i},\tau_{s_{i}}}$ is decisive. Particularly, MTRE-$\tau$ learns two global thresholds $C_{b}<0<C_{u}$ over all $L_{s_{i},\tau_{s_{i}}}$ obtained from the collected OOF pairs ($z^{c}_{s_{i},t}$, $Y_{i}$) that maximize a deployment-aligned metric (Auc, PR-Auc, or $F_{1}$ at a target FPR). Formally, given a hard cap on the number of tokens $T_{\max}\!\leq\!T_{i}$ [^1] In our experiments we cap to the first $T_{i}\leq 10$ tokens for efficiency, handling variable-length sequences and ragged batches is described in App. B.2.1.,the evidence length $\tau_{i}$ for sentence $s_{i}$ induced by $(C_{u},C_{b},T_{\max})$ is
    $\tau_{s_{i}}\;=\;\min\left\{\,t\leq T_{\max}:\;L_{s_{i},\tau_{s_{i}}}\geq C_{u}\;\text{ or }\;L_{s_{i},\tau_{s_{i}}}\leq C_{b}\;\text{ or }\;t=T_{\max}\,\right\}.$ (5)
    Algorithm 4 provides the corresponding pseudo-code to calibrate $\tau_{s_{i}}$ based on the choice of $(C_{u},C_{b},T_{\max})$. Intuitively, when $L_{s_{i},\tau_{s_{i}}}$ is determined to be "decided" by $C_{u}$ or $C_{b}$, there are no further adjustments to the aggregated evidence for the sentence $s_{i}$, otherwise evidence is aggregated up to $T_{max}$.
  4. 4.
    Inferencing with $C$ and $\tau$. Given the calibrated $C^{*}$ and the predicted evidence length $\tau_{s_{i}}$, we can finally conduct inference on the testset $\mathbb{S}_{\text{test}}$ using the Sentence-Level Aggregation Algorithm 4 with $\tau_{i}$ induced by $(C_{u},C_{b},T_{max})$ and $z^{c}_{i,t}$. No thresholds are tuned on $\mathbb{S}_{\text{test}}$.
Throughout this work, we experiment with both MTRE and MTRE-$\tau$, and find that each variant demonstrates effectiveness across multiple tasks. Further details are provided in the experiments section. An alternative stepwise formulation of the MTRE-$\tau$ process (steps 1–4) is presented in Algorithm 5.
5 Experimental results
We begin by outlining the experimental settings adopted in our study. We then present results on MAD-Bench and Safety-Bench, followed by an evaluation on arithmetic-centric tasks and the MathVista benchmark, which are particularly challenging for VLMs and prone to hallucinated outputs. Finally, we report the computational complexity introduced by MTRE during inference and training.
5.1 Experimental setting
Our experimental evaluations are conducted on MathVista (Lu et al., 2023), MM-Safety-Bench (Liu et al., 2023b), MAD-Bench (Li et al., 2023), and four arithmetic/counting tasks from (Rahmanzadehgervi et al., 2024) (see Appendix A). We evaluate outputs from open-source VLMs—LLaVA-v1.5 (7B) (Liu et al., 2023a), mPLUG-Owl (Ye et al., 2023), LLaMA-Adapter V2 (Gao et al., 2023), and MiniGPT-4 (Zhu et al., 2023)—using the 7B versions unless specified. All prompts are listed in Appendix A.1. Prior work shows these models can produce unsafe or unreliable content. We compare MTRE against four baselines: TokenSAR (Duan et al., 2024), Linear Probing (Zhao et al., 2025), Sequential Log-Prob (Guerreiro et al., 2023), and P(True) (Kadavath et al., 2022a) (see Appendix D).
The detections methods are evaluated on two VLM’s types of responses. Type 1 task asks the VLM to directly answer benchmark questions (Direct Answering). Type 2 queries prompt the VLM to evaluate its own outputs (Self Evaluation). For MM-Safety-Bench and MAD-Bench, we additionally follow Zhao et al. (2025) and use three prompt styles: (1) OE, the original open-ended question; (2) OEH, the same question with a hint about possible unanswerability, harmfulness, or deception; and (3) MQ, a meta-question such as “Is this question answerable?”. These prompt variations are used only to diversify outputs for evaluation; each method is applied to responses from either Type 1 or Type 2. We assess MTRE using accuracy, F1, and AUROC. Results are compared to linear probing and P(True), with metrics computed against ground truth. For score-based baselines, we apply the Youden index cutoff Fluss et al. (2005) derived from training scores to compute accuracy and F1 on validation.
5.2 Results on MAD and Safety-Bench
Figure 4 and 5 present the comparative performance of multiple detection methods on the MAD-Bench and MM-Safety-Bench datasets, evaluated under Type I Direct-answering and Type II Self-evaluation tasks, respectively.
As shown in Figure 3, MTRE-based approaches achieve consistently higher performance than baseline methods in Type 1 task. The advantage of MTRE methods is more pronounced on AUROC, where they approach near-perfect discrimination, while baselines demonstrate considerably lower values and larger variances. A similar trend is observed in the F1 score, where MTRE methods dominate and baseline methods lag significantly.
For the Type 2 task, we use VLM output logits across 24 distinct configurations (3 prompts × 4 VLMs × 2 datasets) to evaluate all methods on their ability to determine whether the VLM’s self-assessment is accurate. Once again, MTRE variants outperform all baselines. The gap between MTRE methods and baselines is especially marked in AUROC and F1, underscoring the robustness of MTRE-based detection. Notably, MTRE (LP), which employs a linear probe as the reliability classifier $f_{\theta}$, achieves consistently stronger results than the Linear Probing procedure described in Zhao et al. (2025), which is restricted to training and evaluation on the initial output logit.
MAD-Bench and MM-Safety-Bench Type IRefer to captionRefer to captionRefer to caption
Figure 4: Detection results on Type 1 Direct-answering task in MAD-Bench and $M$-Safety-Bench. (For scores in table format see Appendix E, Tables 89, and 10).
MAD-Bench and MM-Safety-Bench Type IIRefer to captionRefer to captionRefer to caption
Figure 5: Detection results on Type 2 Direct-answering task in MAD-Bench and $M$-Safety-Bench. (For scores in table format, see Appendix E, Tables 56, and 7 ).
5.3 Results on Arithmetic and MathVista
Table 1: Detection performance on Arithmetic and MathVista Type 1 Direct-answering tasks.
Circles Triangles Lines Squares MathVista
Method Acc Auc F1 Acc Auc F1 Acc Auc F1 Acc Auc F1 Acc Auc F1
Lin. Prb. 81.40 86.01 73.50 85.20 84.54 72.72 84.08 88.80 88.61 68.49 56.64 52.03 69.42 70.63 76.81
SAR 59.45 50.10 46.73 65.40 64.46 54.82 53.37 56.21 55.53 41.94 37.84 37.53 54.34 56.48 53.52
Seq Scoring 57.27 46.43 36.02 72.75 67.38 61.74 52.12 51.88 55.92 46.45 49.61 43.71 55.11 56.85 55.82
P(True) 61.80 68.88 58.69 71.95 63.29 67.01 54.45 54.70 38.96 56.61 58.10 37.34 65.86 60.53 53.08
MTRE 87.38 94.38 85.26 90.20 93.66 85.61 87.91 87.79 91.67 97.37 95.68 97.99 76.93 77.54 83.37
MTRE (LP) 85.20 94.69 86.70 86.93 87.20 82.58 79.50 81.13 86.87 87.45 80.79 90.97 76.15 74.29 83.72
MTRE-$\tau$ 85.63 89.50 83.21 89.38 90.48 84.58 85.79 86.12 90.68 96.75 92.11 97.63 76.13 78.35 83.27
MTRE-$\tau$ (LP) 84.81 91.80 86.44 85.29 85.46 80.97 78.87 82.48 86.62 87.20 79.90 90.83 75.92 76.76 83.79
Table 2: Detection performance on MathVista Type 2 Self Evaluation tasks.
LLAVA-7B LLAMA-Adapter MPLUG-Owl MiniGPT-4
Method Acc Auc F1 Acc Auc F1 Acc Auc F1 Acc Auc F1
Lin. Prb. 66.5 70.32 72.00 67.5 71.79 74.60 66.7 72.45 67.27 66.6 70.61 74.16
SAR 56.9 63.10 56.30 63.2 61.30 68.99 53.5 49.04 42.18 63.1 62.53 72.31
Seq Scoring 57.8 62.77 57.75 63.3 61.31 69.04 51.8 49.22 35.36 65.1 62.33 74.49
P(True) 56.6 62.22 48.79 69.0 68.52 43.77 50.9 25.25 00.00 36.4 35.52 53.16
MTRE 78.1 84.40 81.30 76.2 79.94 81.94 75.5 81.02 78.13 74.8 79.30 81.29
MTRE (LP) 76.7 82.85 81.81 75.6 77.40 82.73 76.5 81.68 78.60 74.0 78.64 80.85
MTRE-$\tau$ 76.6 82.91 81.32 75.5 78.12 81.54 74.3 80.02 78.13 74.3 79.77 80.81
MTRE-$\tau$ (LP) 77.1 80.40 82.60 75.0 79.22 82.47 76.1 81.26 78.82 74.5 79.20 81.18
Table 1 reports the performance of various detection methods across four synthetic arithmetic datasets (Circles, Triangles, Lines, Squares) and the MathVista benchmark in Type 1 Direct-answering task (See Tables 111213, and 14 for results sorted by VLM). We observe that baseline methods such as Linear Probe, SAR, Sequence Scoring, and P(True) show mixed results, with performance varying considerably across datasets. For example, Linear Probe achieves relatively strong AUC on the Lines dataset (88.80) but fails to maintain consistent accuracy and F1 on more challenging datasets such as Squares and MathVista. Similarly, SAR and Seq Scoring exhibit limited effectiveness, often trailing behind Linear Probe in both AUC and F1. In contrast, the proposed MTRE family of models consistently outperforms all baselines across nearly all datasets and metrics. MTRE achieves the highest accuracy and F1 on Circles (87.38, 85.26), Triangles (90.20, 85.61), and Squares (97.37, 97.99), while also delivering superior robustness on MathVista (76.93 accuracy, 83.37 F1).
Table 2 further evaluates detection performance on MathVista’s Type 2 self-evaluation tasks. Baseline approaches again show limited performance, with accuracy typically hovering around 556755–67 and F1 values varying unpredictably. By contrast, MTRE demonstrates a clear advantage across all backbones. For LLAVA-7B, MTRE achieves $78.1$ accuracy and $81.3$ F1, substantially outperforming Linear Probe ($66.5$ accuracy, $72.0$ F1). Similarly, on LLAMA-Adapter, MTRE improves accuracy to $76.2$ with a robust 81.9 F1, again exceeding all baseline methods. Comparable gains are observed with MPLUG-Owl and MiniGPT-4, where MTRE and its variants consistently provide improvements in both AUC and F1. Interestingly, the LP and $\tau$ variants of MTRE often yield complementary benefits—for example, MTRE-$\tau$(LP) achieves the best F1 on LLAVA-7B (82.6), while MTRE(LP) provides the strongest overall results on MPLUG-Owl (accuracy = 76.5, F1 = 78.6). Overall, these findings confirm that MTRE is not only effective for direct-answering tasks but also excels in the more nuanced self-evaluation setting, adapting well across multiple model architectures. The results collectively underscore MTRE’s robustness, demonstrating its capacity to provide reliable detection performance in both synthetic and real-world multimodal reasoning benchmarks.
5.4 Computational Cost
Table 3 presents the overhead of MTRE when applied to VLMs. Specifically, MTRE leverages a much smaller model, requiring only about 26.14 MB of VRAM and introducing roughly 4 million additional parameters. This lightweight design offers a substantial efficiency advantage compared to sampling-based approaches (Kuhn et al., 2023), which requires the VLM to generate multiple responses per query for hallucination detection.
Table 3: Computational overhead of MTRE. The inference overhead are averaged among all VLMs.
Metric MTRE reliability classifier Inference overhead
Parameters 4,328,203 $\leq 1\%$
Peak VRAM usage 26.14 MB $\leq 1\%$
Average Inference Time 0.944 ms (per detection) $\leq 1\%$